# Lightweight Joint Optimization of General-Purpose Vision-Language Models and Retrievers for RAG-Based Medical Diagnosis

### v3

@@ Abstract
Retrieving relevant visual and textual information from medical literature and hospital records can enhance diagnostic accuracy for clinical image interpretation. We develop a multimodal retrieval model jointly optimized with an LVLM for medical diagnosis, unlike standard RAG which doesn’t backpropagate LVLM errors to the retriever. Using only general-purpose backbones with lightweight fine-tuning, our model achieves competitive results with medically-pretrained models on clinical classification and VQA tasks. In a novel analysis, we find that different top-retrieved images often yield different predictions for the same target, and that these cases are challenging for all models, even for non-retrieval models. Our joint retrieval optimization significantly improves these cases over standard RAG. However, oracle analysis reveals that while the correct diagnosis is frequently achievable using one of the top retrieved images, in practice there is a large performance gap from the oracle, and rerankers using frontier LVLMs do not close this gap—leaving ample room for improvement by future methods. Code available at https://github.com/Nirmaz/JOMED.
Lightweight Joint Optimization of General-Purpose Vision-Language Models and Retrievers for RAG-Based Medical Diagnosis
@@ 1 Introduction
Figure 1: We jointly optimize a multimodal retriever and a Large Vision-Language Model for medical tasks. We achieve competitive results without resource-intensive medical pre-training and significantly improve performance on challenging cases where different retrieved images lead to inconsistent retrieval prediction.
@@
In this work, we explore a lightweight fine-tuning approach for taking a general-purpose LVLM and optimizing it together with a general-purpose multimodal retriever for medical diagnosis tasks. Different from the common multimodal RAG setup, we perform joint optimization of the LVLM and retriever. This teaches the LVLM to utilize retrieved knowledge—images, captions and clinical reports from both literature and hospital records—while at the same time tuning the retriever to find information that leads to correct predictions by the LVLM. Poorly optimized retrieval mechanisms can mislead models Yoran et al. (2023); Sun et al. (2024); as shown in Figure 2, a RAG model incorrectly diagnoses a benign ultrasound when using a cancerous retrieved image, but joint optimization leads to correct retrieval and prediction.
@@
Our method involves sequential multimodal training with a dual-head retriever architecture and a customized retrieval loss and visual question answering (VQA) training recipe. Our method achieves competitive results in medical image classification and VQA. Importantly, our focus in this work is not to chase SOTA results but to shed light on several interesting observations. One, that a simple and effective joint LVLM-retriever optimization mechanism can lead to a substantial boost in results over current multimodal RAG methods and should thus be more widely adopted for medical diagnosis. Second, that this mechanism achieves competitive results by using models with no medical pre-training (for neither the LVLM nor the retriever) and only lightweight fine-tuning. This resonates with recent findings showing general-purpose LVLMs can rival their medical counterparts Jeong et al. (2024) and has potentially important implications considering the resource-intensive nature of pre-training processes.
@@
Third, previous work on joint optimization in the general domain required large-scale pre-training followed by few-shot task-specific adaptation Izacard et al. (2023); Hu et al. (2023); Lin et al. (2024). In contrast, we conduct joint optimization directly on downstream tasks with no pretraining and only lightweight fine-tuning (as few as 546 samples). This provides first evidence about the feasibility and utility of lightweight data-efficient joint optimization directly on downstream tasks as opposed to in the pre-training setting. We are also the first to explore and show the utility of joint optimization in the multimodal medical domain.
@@
Fourth and importantly, we conduct a novel analysis that finds that our method helps in particular to address a class of errors we term inconsistent retrieval predictions. Inconsistent retrieval predictions are cases in which for a given patient image, each retrieved image leads the model to make different predictions. These cases commonly occur across our experiments and are substantially more difficult for models, both retrieval-augmented and non-retrieval models. This instability with respect to different retrieval candidates leads to high prediction entropy/uncertainty across classes, degrading not only overall RAG results but also their reliability Lambert et al. (2022). As we show, our joint optimization mechanism significantly mitigates this issue and achieves large improvement over standard fine-tuned RAG on these cases. We further discover that for a large proportion of these inconsistent cases, at least one retrieved candidate enables the model to predict the correct answer in an oracle setting in which we provide the model with the correct retrieved candidate. This suggests useful information is being retrieved but is often lost among less helpful candidates. To address this, we explored using the powerful multimodal model o3 OpenAI (2025) as a reranker to identify the most predictive instances. We found it did not close the performance gap to the oracle and was comparable or inferior to our method, leaving significant room for future improvement.
@@
  • We jointly optimize a multimodal retriever and an LVLM for medical classification and VQA. We use only general-purpose backbones without any medical pretraining, and show that by retrieving medical information and co-optimizing the retriever and LVLM we can close much of the gap to expensive medically pre-trained models.
  • Unlike previous general-domain joint optimization that required pre-training, we conduct joint optimization directly on downstream tasks with only lightweight data-efficient fine-tuning.
  • We identify inconsistent retrieval predictions, challenging cases where different retrieved images lead to different predictions for the same query. Our joint optimization method significantly improves performance on these cases when compared to standard RAG.
Figure 2: joint optimization impacts inconsistent retrieval predictions. After joint optimization, retrieved images show greater alignment with query image labels. In Breast, a cancer-free ultrasound query (a) initially retrieved a lesion image (b), however, after joint optimization, the retrieved image is less directly related to the wrong label (c). In VQA-RAD, retrieval shifted from an unrelated medical condition (e) to an image depicting the condition of the query image.
@@ 2 Related Work
Retrieval Augmented joint optimization has been explored primarily in unimodal (text only) NLP with work such as ATLAS Izacard et al. (2023), which conducted large-scale pretraining of both reader and retriever models and evaluated in zero-shot and few-shot scenarios in the general domain. REVEAL Hu et al. (2023) extended this to the multimodal setting, achieving state-of-the-art results in visual tasks in the general domain with extensive pre-training. Shi et al. (2023) proposed training textual retrievers based on reader performance, influencing subsequent joint-optimization methods. Lin et al. (2024) introduced another joint optimization variant and proposed sequential training: first training the reader with retrieval augmentation, then training the retriever. Siriwardhana et al. (2023) demonstrated joint optimization’s effectiveness for domain adaptation. Our work differs by being the first to explore joint optimization optimized directly for downstream tasks, with minimal fine-tuning, as opposed to requiring large-scale pre-training. Our extensive novel analyses are the first to shed light on inconsistent predictions and how joint optimization on downstream tasks helps address them.
Multimodal Retrieval Augmentation in Medical Applications. Multimodal retrieval augmentation began with encoder-based architectures Yuan et al. (2023), which integrated retrieved text and images with query modalities. With Large Vision-Language Models (LVLMs), He et al. (2024a) proposed retrieving relevant labeled examples during inference. Xia et al. (2024b) addressed LVLMs’ misalignment with medical knowledge using retrieval-augmented generation to fetch similar medical reports, employing contrastive learning and context selection strategies. Xia et al. (2024a) proposed domain-aware retrieval for diverse medical data sources, while Wu et al. (2025) combined retrieval augmentation with knowledge graphs.
@@ 3 Method
Figure 3: The two-phase Jomed training. The LVLM is first trained with a frozen retriever on augmented prompts containing retrieved image–caption/report pairs. With the LVLM frozen, the dual-head multimodal retriever is optimized using a KL divergence computed over a selected subset of the model’s logits. An LLM converts open-ended questions into closed form (during training only), which improves results.
Overview. In this section, we present our methodology, termed Jomed (Joint Optimization of Multimodal Retrieval for Enhanced Medical Diagnosis). Our task involves medical image classification and visual question answering. Jomed comprises two main components: a multimodal retriever and a reader. Given an input patient image and a diagnostic question, the multimodal retriever identifies relevant medical knowledge—images and their associated captions or hospital reports—that provide predictive information. The reader then analyzes the retrieved candidates along with the question to generate an answer. We jointly optimize the reader to better leverage the retrieved information and the multimodal retriever to enhance its ability to select informative candidates for the reader. Figure 3 illustrates our framework.
Unlike previous work He et al. (2024a); Xia et al. (2024b, a) our method does not use medical pretraining, and instead we use readily available general LVLMs (Pixtral and Qwen2-vl) and a general retriever, jina-clip Xiao et al. (2024). The selected LVLMs can process multiple images and texts, allowing us to incorporate both the retrieved image and the retrieved caption/report into the model context. We also compare our results to Med-Flamingo Moor et al. (2023). Med-Flamingo is a medical LVLM which is also able to process multiple images, unlike other medical LVLM available at the time of conducting our experiments. We now elaborate on the details.
3.1 Reader Retrieval Augmentation Fine-Tuning
@@
3.2 Sequential Multimodal Retriever Fine-Tuning
Our dual-head multimodal retriever has a text-based head for retrieving relevant (image, text) pairs based on textual similarity to the query, and an image-based head for retrieving pairs based on visual similarity. To train this retriever, we adopt a sequential training strategy, first optimizing the text retrieval head followed by the image retrieval head. During this process, the reader model remains frozen, focusing solely on improving the retriever’s embedding space. Similarly to Izacard et al. (2023) we minimize the KL divergence between the LVLM’s posterior distribution over retrieved pairs and the retriever’s distribution, however we do so directly on downstream tasks with only lightweight fine-tuning in the data-efficient regime without pre-training, and we customize the loss and the training recipe for open-ended questions leading to empirical gains.
Let $z_{k}=(\mathbf{i}_{k},\mathbf{t}_{k})$ denote the $k$-th retrieved image-text pair and $q=(\mathbf{i},\mathbf{q})$ denote the query pair. The posterior distribution reflects the model’s confidence in predicting the correct answer $\mathbf{a}$ given a retrieved pair $z_{k}$ and query $q$:
$p_{k}\propto p_{\text{LVLM}}(\mathbf{a}\mid z_{k}\circ q),$
where $p_{\text{LVLM}}(\mathbf{a}\mid z_{k}\circ q)$ is the probability assigned by the LVLM to $\mathbf{a}$, and $\circ$ denotes prepending the retrieved chunk to the query.
@@ 3.3 Inference
where $p_{\text{R}}(z_{k}\mid q)$ is:
$p_{\text{R}}(z_{k}\mid q)=\frac{\exp(s(q,z_{k}))}{\sum_{j=1}^{K}\exp(s(q,z_{j}))},$
and $s(q,z_{k})$ is the similarity score between the query and the retrieved chunk.
@@
Jomed + PDist .86 .83 .73 .47 .63 .49 .57 .09 .47 .29 .65 .43
Jomed .87 .84 .76 .50 .65 .50 .57 .14 .49 .37 .67 .47
Jomedoracle .87 .84 .85 .64 .69 .51 .64 .14 .52 .41 .71 .51
@@
Jomed + PDist .88 .84 .80 .62 .60 .47 .50 .09 .45 .35 .65 .47
Jomed .90 .87 .80 .62 .60 .51 .56 .14 .51 .37 .67 .50
Jomedoracle .93 .90 .83 .67 .63 .54 .67 .15 .57 .41 .73 .53
@@
Jomed + PDist .77 .45 .89 .81 .92 .32 .86 .53
Jomed .79 .48 .90 .84 .93 .38 .87 .57
Jomedoracle .86 - .92 - .95 - .91 -
@@
Jomed + PDist .75 .41 .89 .81 .89 .32 .84 .51
Jomed .76 .45 .90 .84 .90 .36 .85 .55
Jomedoracle .88 - .92 - .95 - .92 -
Table 1: Results for classification (a) and VQA (b) with non-medically pretrained LVLM backbones (see Tables 3, 4 for comparisons to with medical pre-training; MMed-RAG’s retrieval component is medically pre-trained). joint optimization consistently leads to the best results. For VQA, we evaluate closed questions using the exact match metric and open questions using token recall. Using an oracle reranker (for classification and closed VQA; §4.2) demonstrates that even larger improvements are achievable with Jomed’s top-retrieved images in some datasets.
MMed-RAG
LLaVaMed
variants
MedDr
variants
GSCo
MedVInT
variants
Medical Pre-Training
Size
600K 600K 255K 255K 177K
Table 2: Medical pre-training size for LVLMs under extensive medical pretraining.
@@ 4 Experiments
Jomed Qwen2-vl .87 .76 .14 .37
JomedPixtral .90 .80 .14 .37
@@ 4.1 Experimental Setup
Our datasets include real-world hospital datasets (BRSET Nakayama et al. (2024) and VinDr‑PCXR Pham et al. (2022)) alongside a variety of classification and visual question answering datasets. We focus on a low-resource data-efficient setting (training sets ranging from 546 to 7007 samples in classification, 1790-19,700 in VQA). Medical image annotation is a resource-intensive task, demanding expert annotators whose availability is limited and expensive. Clinical AI also often has poor generalization across heterogeneous medical centers and patient populations, often requiring that models be trained for the specific context in which they are deployed Casey (2022). Such constraints often restrict researchers to datasets comprising only a few thousand samples for a given study. Retrieval augmentation is well-suited for this setting, as it is known to benefit low‑data regimes the most by leveraging external knowledge.
We consider three types of classification tasks: binary, multi-class, and multi-label. These include BreastMNIST (“Breast”; breast ultrasound imaging, binary) Al-Dhabyani et al. (2020), DermaMNIST (“Derma”; pigmented skin lesion images, multi-class) Tschandl et al. (2018), RetinaMNIST (“Retina”; retinal fundus images with multi-class disease annotations) Liu et al. (2022), and two challenging real-world multi-label datasets: VinDr-PCXR (a chest X-ray dataset with 15 labels) Pham et al. (2022) and BRSET (an ophthalmology dataset with 14 labels) Nakayama et al. (2024). For visual question answering, we use widely adopted benchmarks: VQA‑RAD Lau et al. (2018), SLAKE‑English Liu et al. (2021), and PathVQA He et al. (2020). Full details are provided in Appendix A.1.
@@
Baselines. We compare to several RAG baselines. RAD He et al. (2024a), a retrieval-based approach for classification tasks, where the label of the most similar training image is used as context. We integrate RAD with both backbone models (Qwen2-VL and Pixtral). MMed-RAG Xia et al. (2024a), a recent state-of-the-art multimodal RAG framework, using two backbones: LLaVA-Med and LLaVA, both paired with a medically pre-trained retriever (CLIP). We also adopt a standard fine-tuned RAG baseline in which the retriever is fixed and only the reader is fine-tuned, corresponding to the first phase of our approach. Finally, we compare our retriever loss with the base Perplexity Distillation Loss (PDist) Izacard et al. (2023), integrated into Jomed in place of our customized retrieval loss. Full implementation details are in Appendix C. For comparison to medically pre‑trained LVLMs, we include competitive baselines: BiomedGPT Luo et al. (2023); three LLaVA-Med Li et al. (2023) variants; two LVLM variants based on PMC CLIP (MedVInT-TE and MedVInT-TD) Zhang et al. (2023b); and three InternVLChen et al. (2024) variants: MedDr, MedDr + RAD He et al. (2024a), and GSCo He et al. (2024b). Details in Appendix C.

### v4

# LVLM-Aware Multimodal Retrieval for RAG-Based Medical Diagnosis with General-Purpose Models

@@ Abstract
Retrieving visual and textual information from medical literature and hospital records can enhance diagnostic accuracy for clinical image interpretation. However, multimodal retrieval-augmented diagnosis is highly challenging. We explore a lightweight mechanism for enhancing diagnostic performance of retrieval-augmented LVLMs. We train a lightweight LVLM-aware multimodal retriever, such that the retriever learns to return images and texts that guide the LVLM toward correct predictions. In our low-resource setting, we perform only lightweight fine-tuning with small amounts of data, and use only general-purpose backbone models, achieving competitive results in clinical classification and VQA tasks compared to medically pre-trained models with extensive training. In a novel analysis, we highlight a previously unexplored class of errors that we term inconsistent retrieval predictions: cases where different top-retrieved images yield different predictions for the same target. We find that these cases are challenging for all models, even for non-retrieval models, and that our retrieval optimization mechanism significantly improves these cases over standard RAG. However, our analysis also sheds light on gaps in the ability of LVLMs to utilize retrieved information for clinical predictions. Code and models available at: https://github.com/Nirmaz/JOMED.
LVLM-Aware Multimodal Retrieval for RAG-Based Medical Diagnosis with General-Purpose Models
@@ 1 Introduction
Figure 1: We optimize a multimodal retriever and a Large Vision-Language Model for medical tasks. We achieve competitive results without resource-intensive medical pre-training and significantly improve performance on challenging cases where different retrieved images lead to inconsistent retrieval prediction.
@@
In this work, we explore a lightweight fine-tuning approach using a general-purpose LVLM with a general-purpose multimodal retriever for medical diagnosis tasks. Unlike standard multimodal RAG, we train an LVLM-aware multimodal retriever to find knowledge—clinical images, captions, and reports from both medical literature and hospital records—that leads to correct LVLM predictions. Poorly optimized retrieval mechanisms can mislead models Yoran et al. (2023); Sun et al. (2024). As shown in Figure 2, a RAG model incorrectly classifies a benign ultrasound image when conditioned on a cancerous retrieved image, whereas our LVLM-aware retrieval optimization leads to the correct predicted diagnosis.
@@
Our method involves sequential multimodal training with a dual-head retriever architecture and a customized retrieval loss and visual question answering (VQA) training recipe. Our method achieves competitive results in medical image classification and VQA. Importantly, our focus in this work is not to chase SOTA results but to shed light on several interesting observations. One, that a simple and effective LVLM-aware retrieval optimization mechanism can lead to a substantial boost in results over current multimodal RAG methods and should thus be more widely adopted for medical diagnosis; beyond downstream performance, our analysis also shows improved relevance of retrieval candidates. Second, that this mechanism achieves competitive results by using models with no medical pre-training (for neither the LVLM nor the retriever) and only lightweight fine-tuning. This resonates with recent findings showing general-purpose LVLMs can rival their medical counterparts Jeong et al. (2024) and has potentially important implications considering the resource-intensive nature of pre-training processes.
@@
Third, related work in the general domain that tuned both retrieval and generation models required large-scale pre-training followed by few-shot task-specific adaptation Izacard et al. (2023); Hu et al. (2023); Lin et al. (2024). In contrast, we conduct our optimization directly on downstream tasks with no pretraining and only lightweight fine-tuning (as few as 546 samples). This provides first evidence about the feasibility and utility of lightweight data-efficient generation-aware retrieval optimization directly on downstream tasks as opposed to in the pre-training setting. In addition to their reliance on extensive training, another gap in this line of work is that the various techniques developed in the LLM community for tuning retrievers to align with LLMs Izacard et al. (2023); Shi et al. (2023); Lin et al. (2024) have yet to be explored in the multimodal setting with LVLMs.
@@
Fourth and importantly, we conduct a novel analysis that finds that our method helps in particular to address a class of errors we term inconsistent retrieval predictions. Inconsistent retrieval predictions are cases in which for a given patient image, each retrieved image leads the model to make different predictions. These cases commonly occur across our experiments and are substantially more difficult for models, both retrieval-augmented and non-retrieval models. This instability with respect to different retrieval candidates leads to high prediction entropy/uncertainty across classes, degrading not only overall RAG results but also their reliability Lambert et al. (2022). As we show, our retrieval mechanism significantly mitigates this issue and achieves large improvement over standard fine-tuned RAG on these cases. We further discover that for a large proportion of these inconsistent cases, at least one retrieved candidate enables the model to predict the correct answer in an oracle setting in which we provide the model with the correct retrieved candidate. This suggests useful information is being retrieved but is often lost among less helpful candidates. To address this, we explored using the powerful o3 model OpenAI (2025) as a reranker to identify predictive candidates. We found it did not close the performance gap to the oracle and was comparable or inferior to our method, leaving significant room for future improvement.
@@
  • We perform LVLM-aware retrieval optimization for a multimodal retriever and an LVLM on medical classification and VQA tasks (including text generation VQA). We use only general-purpose backbones without medical pretraining, and show that by retrieving relevant medical evidence and training the retriever to guide the LVLM, we close much of the gap to expensive medically pre-trained models and achieve superior performance compared to general-purpose RAG methods.
  • Unlike previous general-domain methods that required pre-training, we conduct lightweight data-efficient optimization of a retriever and LVLM directly on downstream tasks.
  • We identify inconsistent retrieval predictions, challenging cases where retrieved images lead to different predictions for the same query. Our method significantly improves performance on these cases over standard RAG.
Figure 2: LVLM-aware multimodal retrieval finetuning impacts inconsistent retrieval predictions. After LVLM-aware multimodal retrieval finetuning, retrieved images show greater alignment with query image labels. In Breast, a cancer-free ultrasound query (a) initially retrieved a lesion image (b), however, after finetuning, the retrieved image is less directly related to the wrong label (c). In VQA-RAD, retrieval shifted from an unrelated medical condition (e) to an image depicting the condition of the query image.
@@ 2 Related Work
Retrieval-Augmented LLM-Aware Optimization has been explored primarily in unimodal (text only) NLP with work such as ATLAS Izacard et al. (2023), which conducted large-scale pretraining of both reader and retriever models and evaluated in zero-shot and few-shot scenarios in the general domain. REVEAL Hu et al. (2023) extended this to the multimodal setting in the general domain with extensive pre-training; REVEAL used an encoder-decoder architecture with a T5 generator—to our knowledge, since REVEAL no work has explored optimizing multimodal retrievers and generators with causal decoder models and modern VLMs. Shi et al. (2023) proposed training textual retrievers based on reader performance, influencing subsequent retrieval alignment LLM-aware optimization methods. Lin et al. (2024) introduced another LLM-aware optimization variant and proposed first training the reader with retrieval augmentation, then training the retriever. Siriwardhana et al. (2023) demonstrated LLM-aware optimization’s effectiveness for domain adaptation. Our work differs by being the first to explore generator-aware retrieval optimization directly for downstream tasks, with minimal fine-tuning, as opposed to requiring large-scale pre-training. Prior work, such as ATLAS and REVEAL, follows a two-step approach—first conducting large-scale pre-training, then evaluating in few-shot or zero-shot settings, while we use a lightweight one-step approach instead. Our method includes a new loss and new training methodologies. Our novel analyses are the first to shed light on inconsistent predictions and methods to help address them.
Multimodal Retrieval Augmentation in Medical Applications. Multimodal retrieval augmentation began with encoder-based architectures Yuan et al. (2023), which integrated retrieved text and images with query modalities. With LVLMs, He et al. (2024a) proposed retrieving labeled examples during inference. Xia et al. (2024b) improved LVLM RAG with contrastive learning and context selection strategies. Xia et al. (2024a) proposed domain-aware retrieval for diverse medical data, while Wu et al. (2025) combined retrieval augmentation with knowledge graphs. All current multimodal medical RAG methods are not LVLM-aware, and the retriever is trained disjointly from the LVLM.
@@ 3 Method
Figure 3: The two-phase CLARE training. The LVLM is first trained with a frozen retriever on augmented prompts containing retrieved image–caption/report pairs. With the LVLM frozen, the dual-head multimodal retriever is optimized using a KL divergence computed over a selected subset of the model’s logits. An LLM converts open-ended questions into closed form (during training only), which improves results.
Overview. In this section, we present our methodology, termed CLARE (Clinical LVLM-Aware Retrieval). Our task involves medical image classification and visual question answering. CLARE comprises two main components: a multimodal retriever and a reader. Given an input patient image and a diagnostic question, the multimodal retriever identifies relevant medical knowledge—images and their associated captions or hospital reports—that provide predictive information. The reader then analyzes the retrieved candidates along with the question to generate an answer. We optimize the reader to better leverage the retrieved information and the multimodal retriever to enhance its ability to select informative candidates for the reader. Figure 3 illustrates our framework.
Unlike previous work He et al. (2024a); Xia et al. (2024b, a) we do not use medical pretraining, and instead we use readily available general LVLMs (Pixtral and Qwen2-vl) and a general retriever, jina-clip Xiao et al. (2024). The selected LVLMs can process multiple images and texts, allowing us to incorporate both the retrieved image and the retrieved text. We also compare to Med-Flamingo Moor et al. (2023), a medical LVLM which is also able to process multiple images, unlike other medical LVLMs available at the time of conducting our experiments. We now elaborate on the details.
3.1 Reader Fine-Tuning
@@
3.2 Sequential Retriever Fine-Tuning
Our dual-head multimodal retriever has a text-based head for retrieving relevant (image, text) pairs based on textual similarity to the query, and an image-based head for visual similarity. We adopt a sequential training strategy, first optimizing the text retrieval head followed by the image head. During this process, the reader remains frozen, solely on improving retriever’s embedding space. Similarly to Izacard et al. (2023) we minimize the KL divergence between the LVLM’s posterior distribution over retrieved pairs and the retriever’s distribution, however we do so directly on downstream tasks with only lightweight fine-tuning in the data-efficient regime without pre-training, and we customize the loss and the training recipe for open-ended questions leading to empirical gains.
Let $z_{k}=(\mathbf{i}_{k},\mathbf{t}_{k})$ denote the $k$-th retrieved image-text pair and $q=(\mathbf{i},\mathbf{q})$ denote the query pair. The posterior distribution reflects the model’s confidence in predicting the correct answer $\mathbf{a}$ given a retrieved pair $z_{k}$ and query $q$: $p_{k}\propto p_{\text{LVLM}}(\mathbf{a}\mid z_{k}\circ q)$, where $p_{\text{LVLM}}(\mathbf{a}\mid z_{k}\circ q)$ is the probability assigned by the LVLM to $\mathbf{a}$, and $\circ$ denotes prepending the retrieved chunk to the query.
@@ 3.3 Inference
where $p_{\text{R}}(z_{k}\mid q)=\frac{\exp(s(q,z_{k}))}{\sum_{j=1}^{K}\exp(s(q,z_{j}))}$, and $s(q,z_{k})$ is the similarity score between the query and the retrieved chunk.
@@
CLARE + PDist .86 .83 .73 .47 .63 .49 .57 .09 .47 .29 .65 .43
CLARE .87 .84 .76 .50 .65 .50 .57 .14 .49 .37 .67 .47
CLAREoracle .87 .84 .85 .64 .69 .51 .64 .14 .52 .41 .71 .51
@@
CLARE + PDist .88 .84 .80 .62 .60 .47 .50 .09 .45 .35 .65 .47
CLARE .90 .87 .80 .62 .60 .51 .56 .14 .51 .37 .67 .50
CLAREoracle .93 .90 .83 .67 .63 .54 .67 .15 .57 .41 .73 .53
@@
CLARE + PDist .77 .45 .89 .81 .92 .32 .86 .53
CLARE .79 .48 .90 .84 .93 .38 .87 .57
CLAREoracle .86 - .92 - .95 - .91 -
@@
CLARE + PDist .75 .41 .89 .81 .89 .32 .84 .51
CLARE .76 .45 .90 .84 .90 .36 .85 .55
CLAREoracle .88 - .92 - .95 - .92 -
Table 1: Results for classification (a) and VQA (b) with non-medically pretrained LVLM backbones (see Tables 3, 4 for comparisons to medical pre-training; MMed-RAG’s retrieval component is medically pre-trained). CLARE consistently leads to the best results. For VQA, we evaluate closed questions using the exact match metric and open questions using token recall. Using an oracle reranker (for classification and closed VQA; §4.2) demonstrates that even larger improvements are achievable with CLARE’s top-retrieved images in some datasets.
MMed-RAG LLaVaMed variants MedDr variants GSCo MedVInT variants
Medical Pre-Training Size 600K 600K 255K 255K 177K
Table 2: Medical pre-training sizes of existing LVLMs.
@@ 4 Experiments
CLARE Qwen2-vl .87 .76 .14 .37
CLAREPixtral .90 .80 .14 .37
@@ 4.1 Experimental Setup
Our datasets include real-world hospital datasets (BRSET Nakayama et al. (2024) and VinDr‑PCXR Pham et al. (2022)) alongside a variety of classification and VQA datasets. We focus on a low-resource data-efficient setting (training sets ranging from 546 to 7007 samples in classification, 1790-19,700 in VQA). Medical image annotation is a resource-intensive task, demanding expert annotators whose availability is limited and expensive. Clinical AI also often has poor generalization across heterogeneous medical centers and patient populations, often requiring that models be trained for each deployment Casey (2022). Such constraints often restrict researchers to datasets comprising only a few thousand samples for a given study. Retrieval augmentation is well-suited for this setting, as it is known to benefit low‑data regimes the most by leveraging external knowledge.
We consider binary, multi-class, and multi-label classification. These include BreastMNIST (“Breast”; breast ultrasound imaging, binary) Al-Dhabyani et al. (2020), DermaMNIST (“Derma”; pigmented skin lesion images, multi-class) Tschandl et al. (2018), RetinaMNIST (“Retina”; retinal fundus images, multi-class) Liu et al. (2022), and two challenging real-world multi-label datasets: VinDr-PCXR (chest X-rays, 15 labels) Pham et al. (2022) and BRSET (ophthalmology, 14 labels) Nakayama et al. (2024). For VQA, we use widely adopted benchmarks: VQA‑RAD Lau et al. (2018), SLAKE‑English Liu et al. (2021), and PathVQA He et al. (2020). Full details are provided in Appendix A.1.
@@
Baselines. We compare to several RAG baselines. RAD He et al. (2024a), a retrieval-based approach for classification, where the label of the most similar training image is used as context; we use RAD with both Qwen2-VL and Pixtral. MMed-RAG Xia et al. (2024a), a recent state-of-the-art multimodal RAG framework in which the retriever is optimized independently from the LVLM and is not aware of the LVLM during training. We evaluated MMed-RAG using backbones LLaVA-Med and LLaVA, both paired with a medically pre-trained retriever (CLIP). We also adopt a standard fine-tuned RAG baseline in which the retriever is fixed and only the reader is fine-tuned, corresponding to the first phase of our approach. Finally, we compare our retriever loss with the base Perplexity Distillation Loss (PDist) Izacard et al. (2023), integrated into CLARE in place of our customized loss. Implementation details are in Appendix C. To compare to medically pre‑trained LVLMs, we include competitive baselines: BiomedGPT Luo et al. (2023); three LLaVA-Med Li et al. (2023) variants; two LVLM variants based on PMC CLIP (MedVInT-TE and MedVInT-TD) Zhang et al. (2023b); and three InternVLChen et al. (2024) variants: MedDr, MedDr + RAD He et al. (2024a), GSCo He et al. (2024b). Details in Appendix C.